# MSCA-Net: Multi-Scale Context Aggregation Network for Infrared Small Target Detection

### v1

@@ III-A Multi-scale enhanced detection attention module
Given a feature map $X$, we apply layer normalization(LN) to standardize the features, enhancing training stability and accelerating model convergence. After normalization, we employ three convolutions with distinct parameter settings to extract features at specific scales or orientations. The first convolution uses a small kernel to capture local fine-grained details, whereas the second adopts a larger kernel to focus on global contextual information. The third operation incorporates dilated convolutions with varying dilation rates to expand the receptive field, thereby balancing global and local feature representations.
@@
For a given input feature $X$, the whole process of MSEDA can be described as follows:
$N$ $=\text{LN}(X)$ (1)
$f(N)$ $=\text{Conv}_{1}(\text{Conv}_{3}(\text{Conv}_{2}(N)))\otimes\text{Conv}_{2}(N)$
$X$ $=X+\lambda_{1}\text{Conv}_{1}(\text{Conv}_{1}(N)\otimes\text{Conv}_{1}(f_{1}(N),f_{2}(N)))$
$N$ $=\text{LN}(X)$
$X$ $=X+\lambda_{2}\text{Conv}_{1}(\text{Conv}_{1}(N)\otimes\text{Conv}_{1}(\text{Conv}_{2}(N)))$
Output $=\text{Embedding}(X)$
where $\text{LN}(\cdot)$ and $\lambda$ are layer normalization and learnable scaling factors, respectively. $\otimes$ Represents element-by-element multiplication that preserves the dimension, $\oplus$ represents element-by-element addition(all the occurrences of $\otimes$ and $\oplus$ below mean the same thing).
$h_{i}$ $=\text{SA}(Q_{i},K_{i},V_{i},r_{i}),\,1\leq i\leq n$ (2)
$X$ $=\text{Conv}(\text{Concat}[h_{1},\ldots,h_{n}])$
@@
III-B Positional convolutional block attention module
Given the inherent limitations of existing small-target detection methods in integrating global semantics and achieving cross-layer feature fusion, this paper introduces the PCBAM module. Its design is inspired by the cross-modal interaction paradigm proposed in [13] and has been optimized to better suit infrared small-target detection tasks. PCBAM deeply integrates the position attention mechanism (PAM) with the channel-spatial attention module CBAM [41] through an adaptive recalibration strategy and a dual-branch parallel architecture. By employing a dynamic weight allocation mechanism, PCBAM ensures the precise alignment of multi-granularity features, establishing global contextual dependencies along the channel dimension while leveraging spatial attention to guide geometric associations among local features. This ultimately forms a parameter-adaptive feature enhancement pathway. The experimental results demonstrate that PCBAM offers significant advantages in optimizing feature weight distributions and facilitating deep interactions between global and local information. To further explore the specific operational principles of the PCBAM module, we next provide a detailed mathematical analysis of its processing of the input feature $F$.
@@ III-B Positional convolutional block attention module
In Equation (3), $\sigma$ represents the sigmoid activation function, while gap and gmp denote the global average pooling layer and global max pooling layer, respectively. The term mlp refers to a multilayer perceptron composed of two fully connected layers, where the first dense layer contains $C$-th channel units, and the second dense layer consists of $\frac{C}{8}$-th channel units. Subsequently, $F_{C}^{\prime}=F_{C}\otimes F$ is passed into the spatial attention module (SAM), where $\otimes$ represents element-wise matrix multiplication. After processing through the SAM layer, $F_{C}^{\prime}$ is transformed into $F_{S}$.
@@
In Equation (4), $f^{7\times 7}$ denotes a convolutional operation with a kernel size of $7\times 7$ and a dilation rate of 4. DL represents a dense layer, whereas ; signifies the concatenation operation.
Next, we discuss the positional attention module. Given an input feature map ${F}\in{R}^{H\times W\times C}$, the processing steps within this module are as follows. The input passes through a convolutional layer, generating three new feature maps ${B}$, ${Z}$, and ${D}$, each with a dimensionality of ${R}^{H\times W\times C}$. These feature maps are then reshaped into ${R}^{N\times C}$ (where $N=H\times W$). Next, matrix multiplication is performed between the transpose of ${Z}$ and ${B}$, followed by the application of a softmax layer, obtaining a spatial attention feature map ${S}\in{R}^{N\times N}$. ${D}$ is subsequently reshaped back to ${R}^{H\times W\times C}$. Finally, an element-wise summation is applied between ${F}$ and the processed feature map, yielding the output ${F}_{P}\in{R}^{H\times W\times C}$. The entire computation is formulated as shown in Equations (5) and (6).
@@
The term $s_{ji}$ quantifies the influence of position $i$ on position $j$, whereas $\alpha$ represents a scaling parameter initialized to zero and progressively learned and adjusted during training. Ultimately, the overall formulation of the PCBAM module is expressed as follows, with its detailed architecture illustrated in Figure 3.
${F}_{\text{PCBAM}}={F}_{C}^{\prime}+{F}_{S}+{F}_{P}$ (7)
Figure 3: Feature fusion process with MSFA module.
III-C Channel aggregation block module
Existing small-target detection models often struggle with inadequate target feature representation, particularly in complex backgrounds or multi-scale scenarios, where background information overwhelms or diminishes target features. To address this challenge, the Channel Aggregation Block (CAB) [14] is introduced to effectively mitigate the limitations in feature representation. As illustrated in Figure 4, CAB functions primarily by aggregating information across different channels, thereby enhancing the expression of critical features while suppressing redundant features. Specifically, it strengthens channel-wise relationships by assigning higher weights to essential channels, improving the model’s sensitivity to small targets and its robustness against complex backgrounds. Beyond enhancing feature distinguishability, CAB facilitates deep multi-scale information fusion, providing more precise feature representations for small-target detection tasks. Additionally, during the decoding process, CAB reweights key features, directing the model’s focus toward salient target characteristics and improving infrared small-target segmentation accuracy. Moreover, it preserves global contextual information while refining local feature capture, ensuring more reliable support for the final segmentation results. To further explore the specific role of the CAB, the following section presents a mathematical analysis of the processes of the input feature $X$.
For a given input feature $X$, CAB is defined as follows:
$Y=\text{GELU}\left(\text{Conv}_{3\times 3}\left(\text{Conv}_{1\times 1}\left(\text{LN}(X)\right)\right)\right)$ (8)
$CA=Y+\gamma_{C}\otimes\left(Y-\text{GELU}\left(\text{Conv}_{1\times 1}(Y)\right)\right)$ (9)
Specifically, $\text{CAB}(\cdot)$ performs channel information collection and redistribution via the GELU from ${R}^{C\times HW}\to{R}^{1\times HW}$. The channel scaling factor $\gamma_{C}$, initialized to zero, redistributes the channel features through the complementary interaction $Y-\text{GELU}\left(\text{Conv}_{1\times 1}(Y)\right)$. The overall representation of the CAB formula is shown below.
$\text{CAB}(X)=X+\text{Conv}_{1\times 1}(\text{CA}(Y))$ (10)
Figure 4: Dyhead structure diagram.
@@ IV Experimental results
This section validates the effectiveness of the proposed method through extensive experiments. Section 4.1 introduces the public dataset used in this study and describes its partitioning strategy. Section 4.2 details the training and testing procedures, including hyperparameter settings and hardware configurations. Section 4.3 presents the evaluation metrics used to ensure the scientific rigor and reliability of the comparative analysis. Section 4.4 conducts a comprehensive comparison between the proposed method and other state-of-the-art infrared small target detection networks from both qualitative and quantitative perspectives to assess the performance advantages. Section 4.5 systematically analyzes the role and contribution of each module within the framework through ablation studies, further verifying their critical impact on overall performance enhancement. Section 4.6 presents experimental results that intuitively illustrate the detection effectiveness of different methods, strongly supporting the experimental conclusions.
@@ IV-A Dataset preparation
To evaluate the performance of the proposed model in infrared small target detection, we conducted experiments on three public data sets: NUAA-SIRST [28], NUDT-SIRST [31] and IRSTD-1K [30]. Specifically, NUAA-SIRST contains 427 infrared images, NUDT-SIRST includes 1,327 images, and IRSTD-1K contains 1,001 images. All three datasets were divided into the training and test sets at a 7:3 ratio. These datasets cover infrared images captured at short, medium, and 950nm wavelengths, covering a wide range of remote sensing backgrounds, including the sky, ground, buildings, and oceans. In addition, the dataset contains multiple target types, including military targets such as drones and ships, further enriching the diversity of test scenarios. At the same time, these datasets also contain many challenging scenes, such as low contrast and complex backgrounds, ensuring a comprehensive assessment of model detection capabilities.
@@ IV-B Implementation details
To ensure the rigor of the comparative experiment, we adopted unified experimental datasets. First, all deep learning models were retained based on the training data set used by MSCA-Net to ensure the consistency of data division and training samples; Secondly, in the model reasoning stage, the traditional algorithm strictly follows the parameter configuration suggested in the original text. For the method requiring threshold segmentation, we fixed the threshold parameters recommended in various literature to avoid the influence of artificial adjustment on the results. https://github.com/XinyiYing/BasicIRSTD provides open-source implementations of most techniques.
In terms of experimental details, all the experiments were conducted on an Intel Xeon Platinum 8481 CPU with a Linux operating system and 24.0 GB of RAM. All the deep learning models were implemented via the torch framework. To accelerate training, an NVIDIA GeForce GTX 4090 D GPU with CUDA 11.8 was utilized. The Adam optimizer with a smoothing constant of 0.9 was employed to optimize the MSE loss function. The initial learning rate was set to $10^{-3}$, with a minimum value of $10^{-5}$. We trained the model for 1,000 epochs, automatically adjusting the learning rate using a cosine annealing strategy during training. The batch size is set to 16, and the IoU threshold for NMS is set to 0.5.
@@ IV-C Evaluation metrics
Segmentation methods based on U-Net primarily employ pixel-level evaluation metrics, including the IoU, nIoU, $P_{d}$, and $F_{a}$. These measures focus on assessing the shape characteristics of the detected targets.
(1)IoU: The IoU (intersection over union) is a pixel-level evaluation metric designed to measure an algorithm’s ability to delineate target contours. It quantifies model performance by computing the ratio between the intersection and union of the predicted and ground truth regions. The definition is as follows:
@@
(2)nIoU: nIoU is a standardized version of the IoU specifically designed for infrared small target segmentation. Its primary objective is to provide a more precise assessment of small target segmentation performance while mitigating the influence of large targets on the results.
@@
(3)$P_{d}$: $P_{d}$ is defined as the ratio of correctly predicted targets $N_{\text{pred}}$ to the total number of targets $N_{\text{all}}$, expressed as:
@@
(4)$F_{a}$ is defined as the ratio of falsely predicted target pixels $N_{\text{false}}$ to the total number of pixels in the image $P_{\text{all}}$, expressed as:
@@ IV-D Detection performance comparison
To validate the superiority of our method, we compared it with several advanced approaches and conducted experiments on three publicly available datasets: NUAA-SIRST, NUDT-SIRST, and IRSTD-1K. For the NUAA-SIRST, NUDT-SIRST, and IRSTD-1K datasets, the mIoU reached 78.43%, 94.56%, and 67.08%, respectively, whereas the nIoU reached 79.91%, 94.36%, and 67.15%, respectively. Table I shows the detailed quantitative results.
TABLE I: Comparison with other prior methods
@@
mIoU nIoU/$P_{d}$ $F_{a}$ mIoU nIoU/$P_{d}$ $F_{a}$ mIoU nIoU/$P_{d}$ $F_{a}$
@@
The experimental results demonstrate that model-driven approaches generally perform worse than data-driven methods do, particularly with a notable gap in terms of the IoU. This is because model-driven methods focus on the overall position of the target rather than precise shape segmentation. However, in complex and dynamic backgrounds, model-driven methods often overlook small targets and their boundaries, leading to poor performance in small target detection tasks.
@@
Figure 5: ROC curves of different methods on the NUAA-SIRST, NUDT-SIRST, and IRSTD-1K datasets. (a) NUAA-SIRST. (b) NUST-SIRST. (c) IRSTD-1K.

### v2

@@ III-A Multi-scale enhanced dilated attention module
Given a feature map $X$, we apply layer normalization (LN) to standardize the features, enhancing training stability and accelerating model convergence. After normalization, we employ three convolutions with distinct parameter settings to extract features at specific scales or orientations. The first convolution uses a small kernel to capture local fine-grained details, whereas the second adopts a larger kernel to focus on global contextual information. The third operation incorporates dilated convolutions with varying dilation rates to enlarge the receptive field, thereby balancing global and local feature representations.
@@
For a given input feature $X$, the whole process of MSEDA can be formally described by the following equations (LABEL:formula:mseda1), (2) and (3):
$Y$ $=X+\lambda_{1}\left(\mathrm{MSLK}\left({LN}(X)\right)\right)$ (1)
$Z$ $=Y+\lambda_{2}\left(\mathrm{EU}\left({LN}(Y)\right)\right)$
where $\mathrm{LN}(\cdot)$ and $\lambda_{i}$ are the layer normalization and learnable scaling factors, respectively. $\otimes$ denotes element-wise multiplication, and $\oplus$ denotes element-wise addition.
Subsequently, the processed features are passed through an embedding layer and then fed into the PConv module, which expands the channel dimension by a factor of three. The output is then reshaped and permuted to produce multiple groups of dilated heads, which are split into the corresponding query, key, and value tensors, denoted as $Q_{i},K_{i},V_{i}$. Formally, this process can be expressed as:
$Q_{i},K_{i},V_{i}=\text{split}\left(\text{res}\left(\text{PConv}(\text{Eb}(Z))\right)\right),\,1\leq i\leq 3\\ \$ (2)
here, $\text{res}(\cdot)$ denotes the reshape operation and $\text{Eb}(\cdot)$ represents the embedding function.
These components are subsequently passed into the Dilated Attention (DA) module for further refinement, as formulated in equation (3):
$h_{i}$ $=\text{SA}(Q_{i},K_{i},V_{i},r_{i}),\,1\leq i\leq 3$ (3)
$O$ $=\text{Conv}_{1\times 1}(\text{Concat}[h_{1},\ldots,h_{n}])$
@@
III-B Positional convolutional block attention
Considering the intrinsic constraints of current small-target detection approaches in integrating global semantics and achieving cross-layer feature fusion, this paper introduces the PCBAM module. Its design is inspired by the cross-modal interaction paradigm proposed in [13] and has been optimized to better suit infrared small-target detection tasks. PCBAM deeply integrates the position attention mechanism (PAM) with the channel-spatial attention module CBAM [42] through an adaptive recalibration strategy and a dual-branch parallel architecture. By employing a dynamic weight allocation mechanism, PCBAM ensures the precise alignment of multi-granularity features, establishing global contextual dependencies along the channel dimension while leveraging spatial attention to guide geometric associations among local features. This ultimately forms a parameter-adaptive feature enhancement pathway. The experimental findings indicate that PCBAM provides notable improvements in refining feature weight allocation and promoting comprehensive interactions between holistic context and fine-grained details. To further explore the specific operational principles of the PCBAM module, we next provide a detailed mathematical analysis of its processing of the input feature $F$.
@@ III-B Positional convolutional block attention
in equation (4), $\sigma$ denotes the sigmoid activation function, whereas gap and gmp denote the global average pooling and global max pooling operations, respectively. The term mlp refers to a multilayer perceptron composed of two fully connected layers, where the first dense layer contains $C$-th channel units, and the second dense layer consists of $\frac{C}{8}$-th channel units. Subsequently, $F_{C}^{\prime}=F_{C}\otimes F$ is passed into the spatial attention module (SAM), where $\otimes$ represents element-wise matrix multiplication. After processing through the SAM layer, $F_{C}^{\prime}$ is transformed into $F_{S}$:
@@
in equation (5), $f^{7\times 7}$ indicates a convolutional operation with a $7\times 7$ kernel and a dilation rate of 4. DL refers to a dense layer, whereas ; represents the concatenation operation.
Next, we discuss the positional attention module. Given an input feature map ${F}\in{R}^{H\times W\times C}$, the processing steps within this module are as follows. The input is processed by a convolutional layer, generating three new feature maps ${B}$, ${Z}$, and ${D}$, each with a dimensionality of ${R}^{H\times W\times C}$. These feature maps are then reshaped into ${R}^{N\times C}$ (where $N=H\times W$). Next, matrix multiplication is conducted between the transpose of ${Z}$ and ${B}$, and a softmax function is applied to yield the spatial attention map ${S}\in{R}^{N\times N}$. ${D}$ is subsequently reshaped back to ${R}^{H\times W\times C}$. Finally, an element-wise addition is applied between ${F}$ and the processed output, yielding the output ${F}_{P}\in{R}^{H\times W\times C}$. The entire computation is formulated as shown in equations (6) and (7):
@@
the term $s_{ji}$ quantifies the influence of position $i$ on position $j$, whereas $\alpha$ represents a scaling parameter initialized to zero and progressively learned and adjusted during training. Ultimately, the complete formulation of the PCBAM module is given by the following expression, with its detailed architecture illustrated in Figure 3.
${F}_{\text{PCBAM}}={F}_{C}^{\prime}+{F}_{S}+{F}_{P}.$ (8)
Figure 3: PCBAM module structure diagram.
III-C Channel aggregation feature fusion block
Existing small-target detection models often struggle with inadequate target feature representation, particularly in complex backgrounds or multi-scale scenarios, where background information overwhelms or diminishes target features. To tackle this issue, the Channel Aggregation Feature Fusion Block (CAB) [14] is introduced to effectively mitigate the limitations in feature representation. As illustrated in Figure 4, CAB functions primarily by fusing the feature information of low-dimensional and high-dimensional, thereby enhancing the expression of critical features while suppressing redundant features. Specifically, it strengthens channel-wise relationships by assigning higher weights to essential channels, improving the model’s sensitivity to small targets and its robustness against complex backgrounds. Beyond enhancing feature distinguishability, CAB facilitates deep multi-scale information fusion, providing more precise feature representations for small-target detection tasks. Additionally, during the features fuse process, CAB performs weighted processing on the key channels, directing the model’s focus toward salient target characteristics and improving infrared small-target segmentation accuracy. Moreover, it preserves global contextual information while refining local feature capture, ensuring more reliable support for the final segmentation results. To further explore the specific role of the CAB, the following section presents a mathematical analysis of the processes of the input feature $F_{1}$ and $F_{2}$.
For the input feature $F_{1}$, its transformation is defined by equation (9):
$F_{1}$ $=\text{Gelu}\left(\text{Conv}_{3\times 3}\left(\text{Conv}_{1\times 1}\left((F_{1})\right)\right)\right)$ (9)
$F_{1}$ $=\text{Dropout}(F_{1})$
similarly, the operation applied to $F_{2}$, its transformation is defined by equation (10):
$\mathrm{SE}$ $=\sigma\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Gelu}\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{pool}(F_{1})\right)\right)\right)\right)$ (10)
$F_{2}$ $=\text{Gelu}\left(\text{SE}\left(\text{Conv}_{1\times 1}\left((F_{1})\right)\right)\right)$
$F_{2}$ $=\text{Dropout}(F_{2})$
specifically, $\mathrm{pool}$ performs average pooling, and $\sigma$ denotes the Sigmoid function. The SE module applies weighted processing to deep features. Then, $F_{1}$ and $F_{2}$ are fused by element-wise average (AVG), followed by element-wise multiplication between the Sigmoid and Dropout-processed $F_{1}$ and the averaged features. Finally, the output is obtained through a Gelu activation. As shown in equation (11):
$\mathrm{Output}=\mathrm{Gelu}\left(\sigma\left(\mathrm{AVG}(F_{1},F_{2})\right)\otimes F_{1}\right).$ (11)
Figure 4: CAB structure diagram.
@@ IV Experimental results
This section validates the efficacy of the proposed method through extensive experiments. Section 4.1 introduces the public dataset used in this study and describes its partitioning strategy. Section 4.2 details the training and testing procedures, including hyperparameter settings and hardware configurations. Section 4.3 presents the evaluation metrics used to ensure the scientific rigor and reliability of the comparative analysis. Section 4.4 conducts a comprehensive comparison between the proposed method and other state-of-the-art (SOTA) infrared small target detection networks from both qualitative and quantitative perspectives to assess the performance advantages. Section 4.5 systematically analyzes the role and contribution of each module within the framework through ablation studies, further verifying their critical impact on overall performance enhancement. Section 4.6 presents experimental results that intuitively illustrate the detection effectiveness of different methods, strongly supporting the experimental conclusions.
@@ IV-A Dataset preparation
To assess the efficacy of our introduced framework in IRSTD, we conducted experiments on three public datasets: NUAA-SIRST [28], NUDT-SIRST [31] and IRSTD-1K [30]. Specifically, NUAA-SIRST contains 427 infrared images, NUDT-SIRST includes 1,327 images, and IRSTD-1K contains 1,001 images. All three datasets were divided into the training and test sets at a 7:3 ratio. These datasets cover infrared images captured at short, medium, and 950nm wavelengths, covering a wide range of remote sensing backgrounds, including the sky, ground, buildings, and oceans. In addition, the dataset contains multiple target types, including military targets such as drones and ships, further enriching the diversity of test scenarios. At the same time, these datasets also contain many challenging scenes, such as low contrast and complex backgrounds, ensuring a comprehensive assessment of model detection capabilities.
@@ IV-B Implementation details
To systematically evaluate the performance of the proposed methods, we compared MSCA-Net with SOTA IRSTD methods. The experimental benchmark consists of two paradigms: three model-driven approaches, Top-Hat [19] based on morphological filtering, WSLCM [22] based on local contrast, and IPI [18] based on low-rank sparse decomposition. In deep learning, we compared seven cutting-edge models: the asymmetric context modulation network (ACM) [28], the attentional local contrast network (ALCNet) [29], the receptive field and direction-induced attention network (RDIAN) [33], the improved U-Net architecture (ISTDU-Net) [34], the interior attention-aware network (IAANet) [43], the dense nested attention network (DNA-Net) [31], and the nested U-Net network (UIU-Net) [32].
To maintain the rigor of the comparative analysis, we employed standardized experimental datasets. First, all deep learning models were retained based on the training data set used by MSCA-Net to ensure the consistency of data division and training samples; Secondly, in the model reasoning stage, the traditional algorithm strictly follows the parameter configuration suggested in the original text. For the method requiring threshold segmentation, we fixed the threshold parameters recommended in various literature to avoid the influence of artificial adjustment on the results. https://github.com/XinyiYing/BasicIRSTD provides open-source implementations of most techniques.
Regarding the experimental setup, all tests were performed on an Intel Xeon Platinum 8481 CPU running a Linux operating system with 24.0 GB of RAM. All deep learning models were developed using the torch framework. To expedite the training process, an NVIDIA GeForce GTX 4090 D GPU with CUDA 11.8 was employed. The Adam optimizer with a smoothing constant of 0.9 was employed to optimize the MSE loss function. The initial learning rate was set to $10^{-3}$, with a minimum value of $10^{-5}$. We trained the model for 1,000 epochs, automatically adjusting the learning rate using a cosine annealing strategy during training. The batch size is set to 16, and the IoU threshold for NMS is set to 0.5.
@@ IV-C Evaluation metrics
Segmentation approaches predominantly utilize pixel-wise evaluation metrics, such as IoU, nIoU, $P_{d}$, and $F_{a}$. These measures focus on assessing the shape characteristics of the detected targets.
(1) IoU: IoU is a pixel-level evaluation metric designed to measure an algorithm’s ability to delineate target contours. It quantifies model performance by computing the ratio between the intersection and union of the predicted and ground truth regions. The definition is as follows:
@@
(2) nIoU: nIoU is a normalized variant of IoU. Its main goal is to provide a more precise assessment of small target segmentation performance while reducing the impact of larger targets on the outcomes.
@@
(3) $P_{d}$: $P_{d}$ is defined as the proportion of accurately detected targets $N_{\text{pred}}$ to the total number of targets $N_{\text{all}}$, expressed as:
@@
(4) $F_{a}$ is defined as the fraction of falsely detected target pixels $N_{\text{false}}$ over the total number of pixels in the image $P_{\text{all}}$, expressed as:
@@ IV-D Detection performance comparison
To demonstrate the effectiveness of our approach,we conducted a comparative study against several SOTA methods using three publicly available datasets: NUAA-SIRST, NUDT-SIRST, and IRSTD-1K. For these datasets, the mIoU reached 78.43%, 94.56%, and 67.08%, respectively, whereas the nIoU reached 79.91%, 94.36%, and 67.15%, respectively. The detailed quantitative results are summarized in Table I.
TABLE I: Comparison with other SOTA methods on NUAA-SIRST, NUDT-SIRST and IRSTD-1K in $\mathrm{mIoU}~{}(\%)$, $\mathrm{nIoU}~{}(\%)$, $\mathrm{P}_{\mathrm{d}}~{}(\%)$, $\mathrm{F}_{\mathrm{a}}(10^{-6})$.
@@
mIoU nIoU $P_{d}$ $F_{a}$ mIoU nIoU $P_{d}$ $F_{a}$ mIoU nIoU $P_{d}$ $F_{a}$
@@
TABLE II: Performance comparison with SOTA methods on IRSTD-1K in terms of parameters(M), Flops(G), inference time(s), and memory usage(M).
Method Params Flops Inference Memory
ACM [28] 0.398 0.402 0.029 897
RDIAN [33] 0.216 3.718 0.035 3331
ISTDU-Net [34] 2.751 7.944 0.039 10717
UIU-Net [32] 50.540 54.425 0.060 14747
DNA-Net [31] 4.696 14.261 0.050 9869
SCTransNet [35] 11.19 20.24 0.067 7077
MSCA-Net(Ours) 9.85 11.79 0.020 11031
The experimental results demonstrate that model-driven approaches generally perform worse than data-driven methods do, particularly with a notable gap in terms of the IoU. This is attributed to the fact that model-driven approaches emphasize the approximate localization of targets rather than achieving accurate delineation of their shapes. However, in complex and dynamic backgrounds, model-driven methods often overlook small targets and their boundaries, leading to poor performance in small target detection tasks.
Table II presents a comprehensive comparison between MSCA-Net and SOTA methods on the IRSTD-1K datasets, covering model size, GFLOPs, inference time for each image, and GPU memory usage. Compared to SCTransNet, MSCA-Net reduces the number of parameters by 11.97% and FLOPs by 41.74%. Notably, despite a 55.87% increase in GPU memory consumption, the inference speed per image is improved by a factor of 3.35, demonstrating a favorable trade-off between processing efficiency and memory usage.
@@
Figure 5: ROC curves of different methods. (a) NUAA-SIRST. (b) NUST-SIRST. (c) IRSTD-1K.